OpenStudy (anonymous):

giving medals! If the midpoint of two points is (5,-1) and one of the points is (4,-3), find the other point.

OpenStudy (anonymous):

@aum

OpenStudy (johnweldon1993):

Well we know that the midpoint is found using \[\large (\frac{x_2 + x_1}{2}, \frac{y_2 + y_1}{2})\]

OpenStudy (johnweldon1993):

We know the 'x' will come out to 5 and the y will come out to -1 so \[\large \frac{x_2 + 4}{2} = 5\] \[\large \frac{y_2 -3}{2} = -1\]

OpenStudy (johnweldon1993):

So you can solve for x2 and y2

OpenStudy (anonymous):

whats x2 and y2

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ -3}})\quad &({\color{red}{ a}}\quad ,&{\color{blue}{ b}}) \end{array}\qquad \left(\cfrac{{\color{red}{ a}} + {\color{red}{ 4}}}{2}\quad ,\quad \cfrac{{\color{blue}{ b}} + {\color{blue}{ (-3)}}}{2} \right)=(5,-1) \\ \quad \\ thus \qquad \cfrac{{\color{red}{ a}} + {\color{red}{ 4}}}{2}=5\qquad and\qquad \qquad \cfrac{{\color{blue}{ b}} + {\color{blue}{ (-3)}}}{2}=-1\)

OpenStudy (johnweldon1993):

That would be what you are solving for x2 would be the x-coordinate of your point and y2 will be the y-coordinate of the point... so for example for the 'x'...we need to solve \[\large \frac{x_2 + 4}{2} = 5\] what does x2 equal?

OpenStudy (anonymous):

6

OpenStudy (johnweldon1993):

Correct, that will be your 'x'....now what about 'y'?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

(6,1)

OpenStudy (johnweldon1993):

Correct again! So your point will be \(\large (6,1)\)

OpenStudy (anonymous):

can you help me with one more

OpenStudy (anonymous):

Find the distance between the following pairs of points: (-3, 6), (3, 2)

OpenStudy (johnweldon1993):

This we can just use the distance formula \[\large d = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\]

OpenStudy (anonymous):

so whats x2 and x1 and y2 and y1=

OpenStudy (johnweldon1993):

Oh I see the confusion now... You have 2 points... Point 1 = (-3 , 6) (x1 , y1) Point 2 = (3 , 2) (x2 , y2) the 1 and the 2 just represent which point the numbers are from

OpenStudy (anonymous):

you mean 5,-1 and 4,-3????

OpenStudy (johnweldon1993):

"Find the distance between the following pairs of points: (-3, 6), (3, 2)"

OpenStudy (anonymous):

ohh sorr confusing from the last problem lol

OpenStudy (anonymous):

sorry

OpenStudy (johnweldon1993):

Lol no problem :)

OpenStudy (anonymous):

so hold on would this be the answer

OpenStudy (anonymous):

52?

OpenStudy (johnweldon1993):

Close but you're forgetting the square root again \[\large d = \sqrt{(3 - (-3))^2 + (2 - 6)^2} = \sqrt{(6^2) + (-4)^2} = \sqrt{36 + 16} = \sqrt{52}\]

OpenStudy (anonymous):

37?

OpenStudy (johnweldon1993):

....?

OpenStudy (anonymous):

36+1

OpenStudy (anonymous):

?

OpenStudy (johnweldon1993):

I'm confused on what you are getting? where are you getting the 36 and the 1?

OpenStudy (anonymous):

you wrote it

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

sq root of 36+1

OpenStudy (johnweldon1993):

\[\large d = \sqrt{(3 - (-3))^2 + (2 - 6)^2} = \sqrt{(6^2) + (-4)^2} = \sqrt{36 + 16} = \sqrt{52}\]

OpenStudy (johnweldon1993):

Oh wait, is it getting cut off for you? \[\large d = \sqrt{(6^2) + (-4)^2} = \sqrt{36 + 16} = \sqrt{52}\]

OpenStudy (anonymous):

7.21 but what the sq root form

OpenStudy (anonymous):

2sq root 13?

OpenStudy (johnweldon1993):

The original distance formula is \[\large d = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\]

OpenStudy (anonymous):

so im right

OpenStudy (johnweldon1993):

\(\large 2\sqrt{13}\) is indeed correct

OpenStudy (anonymous):

ty!

OpenStudy (johnweldon1993):

Of course!